Status Report

NASA ISS On-Orbit Status 19 June 2013

By SpaceRef Editor
June 19, 2013
Filed under , , ,
NASA ISS On-Orbit Status 19 June 2013
NASA ISS On-Orbit Status 19 June 2013

International Space Station (ISS) Management Center (IMC) Daily Summary Report
Increment 36 Operations – Stage 36-6
06/19/13 4:00 pm Central Daylight Time (CDT)
170/21:00 Greenwich Mean Time (GMT)

New Status: Biological Rhythms 48-hour Operations: Parmitano attached electrodes and will wear a Digital Walk Holter Electrocardiogram (ECG) for 48 hours. The investigation examines the circadian variations in the cardiac autonomic function during space flight. Results will be used to improve crew health technology on long-duration space flights.

FASES Experiment: Parmitano installed the FASES experiment container in the Fluids Science Laboratory (FSL) and performed the first run of this experiment. The forty-four samples of the FASES studies emulsion coarsening kinetics in microgravity to investigate the link between emulsion stability and the physical-chemistry of droplet interfaces. Results will be used to generate a model of emulsion dynamics which can be used in industrial applications.

EVA Flight Procedures Study: Parmitano, Misurkin and Yurchikhin performed a flight procedures study in preparation for Russian EVA #33 scheduled for Monday, June 24.

ATV4 Operations: All 3 USOS crew members continued transferring cargo from ATV4 to the ISS. The crew completed the cargo operations planned for the first docked week – an estimated 19 hours worth of cargo operations completed in approximately 10 hours. Cargo specialists will be providing additional ATV4 unload plans tomorrow for the crew to begin executing.

ATV4 Reboost: ATV4 performed a reboost using the Orbital Control Systems (OCS) 2 and 4 thrusters for 6 minute 47 seconds, increasing the ISS velocity by 0.98 m/s and the altitude by 1.73 km. This was the first maneuver for the joint ATV4-ISS vehicle. The purpose of the reboost was to test the ATV thrusters and to set up phasing for 52P launch on July 27 for 4-orbit rendezvous.

51 Progress Deorbit: 51P performed six burns in support of the non-ISS-related “Radar-Progress-8” geophysical experiment (8th in the Radar-Progress series) on June 13 – 18, 2013. All burns were nominal. Progress performed a nominal 101.356 m/s deorbit burn at 7:53 am CDT.

Advanced Resistive Exercise Device (ARED) Cable Arm Ropes Re-installation – Cassidy removed and re-installed the Cable Arm Ropes on the ARED to address concerns with the Cable Arms not staying engaged in their detents during bar exercises. Following the re-installation, the Cable Arms remained engaged in the detents during the subsequent exercise session. The crew is go to use ARED for all exercises.

Mobile Servicing System (MSS) Operations: On Tuesday night, Robotics Ground Controllers powered up the MSS and maneuvered the Space Station Remote Manipulator System (SSRMS) to stow the Special Purpose Dexterous Manipulator System (SPDM) on Mobile Base System (MBS) Power Data Grapple Fixture (PDGF) 2. They then maneuvered the SSRMS to the start position for the Free Flyer Release On-Orbit Checkout Requirement (OCR). Ground Controllers then performed the OCR, the results of which will be used to evaluate the improvement in SSRMS performance during free flyer release pullbacks using the Friction Feed Forward (FFF) Compensation files uplinked earlier today. At the completion of the OCR, the SSRMS was maneuvered to a park position clear of the Space Communication and Navigation (SCAN) Stay Out Zone.

Today’s Planned Tasks
– PROK. Urine РН Test [Complete]
– SM ПСС (Caution & Warning Panel) Test [Complete]
– Laptop RS2(1) Reboot / Onboard Computer System (БВС) [Complete]
– Daily Planning Conference (S-band). [Complete]
– Regeneration of БМП Ф1 Micropurification Cartridge (start) [Complete]
– СОЖ Maintenance
– Removal of two Handles from ТКГ Hatch Front [Complete]
– LCS. SW Installation and БТЛС Data Download from БСПН to RSS1 laptop. r/g 2678 [Complete]
– Removal of two handles from the front of ТГК hatch, [Complete]
– Columbus Front Deck Rack Front Restow [Complete]
– Replacement of urine receptacle (МП) and filter-insert r/g 2679 [Complete]
– ATV Cargo Operations [Complete]
– Space Suit Comm and Telemetry Config, Tranzit-B Activation [Complete]
– USOS Window Shutter Close [Complete]
– Beta-08 Data Output Device (УСИ) Checkout Using Gamma-1M r/g 2674 [Complete]
– Orlan and БСС Telemetry Check. Spacesuit Voice Comm Check And Biomed [Complete]
– Comm reconfig after Orlan Comm. Check / EVA from DC, [Complete]
– Review airlock ops procedure per RODF “EVA from DC1”. r/g 2672 [Complete]
– On MCC Go ISS Repress from ТКГ 418 СрПК Section 2 (DC1) (start) [Complete]
– PROK. Diet Logging [Complete]
– ISS O2 Repress from ТКГ 418 (DC2) СРПК Section 2 (terminate) [Complete]
– EMU LIB battery Charge (end) [Complete]
– ARED rope replace [Complete]
– Setting up NASA gear for use on Orlan-M [Complete]
– High definition Video Camera Assembly Activation for High Definition Mode [Complete]
– Photo of Traces Left by ATV СтМ Probe on SM АСП-О after ATV4 Docking / r/g 2664 [Complete]
– CIR – Hardware Teardown [Complete]
– Installation of US add-on hardware on Orlan-M Photography of Orlan Gear / r/g 2673 [Complete]
– PAO Hardware Setup [Complete]
– Docking Cone Internal Surface Photos/Video Downlink [Complete]
– Crew Prep for PAO [Complete]
– PAO Event [Complete]
– SM Interior Reconfig to Initial Position after Removal of PCE Box / r/g 2646 [Complete]
– BLR48 – Starting measurements [Complete]
– ATV Cargo Operations [Complete]
– Cleaning ventilation screens on FGB interior panels [Complete]
– Installation of USOS EMU Lights and Video Camera On Orlan-MK No. 6, No. 5 [Complete]
– IDENTIFIKATSIYA. Copy ИМУ-Ц micro-accelerometer data to laptop / r/g 1813 [Complete]
– ATV – Transfers Tagup prep [Complete]
– RGN – Initiate drain into EDV [Complete]
– HRF. Urine Sample Insertion [Complete]
– Pro K Urine PH Kit Unstow [Complete]
– ISERV – Opening window shutters in the LAB [Complete]
– RGN – Terminate drain into EDV [Complete]
– MO-21 Prep. ECOSPHERA Battery Charge r/g 2641
– CRYOGEM-03 refrigerator Setup [Complete]
– HAM Radio Session with International Student Camp “Slavic Commonwealth r/g 2660 [Complete]
– IMS Update
– ATV Transfers Tagup [Complete]
– Daily Planning Conference (S-band) [Complete]
– БМП Ф1 Absorption Cartridge Regeneration (termination) [Complete]

Task List
– CEO ops [Complete]
– WHC KTO replace [Complete]
– Preparation of reports for Roskosmos site
– ECON-M. Observations and Photography
– URAGAN Observations and Photography
– ESA record video msg [Complete]
– CMS BD2 harness check
– EVA-Connector sleeve locate [In work]

– Nominal Commanding

Three-Day Look Ahead:
Thursday, June 20: ATV4 closeout imagery/cargo transfer, BCAT, Spinal, Ultrasound
Friday, 6/28: ATV cargo transfer, Alloy, Radin, Robonaut, Space Headaches
Saturday, 6/29: Crew off duty

SpaceRef staff editor.